OpenStudy (raylynn):

X^2-2x-15 X-5 ___________ / _________ X-3 X^2-9

zepdrix (zepdrix):

Hello Ms Raylynn!\[\large\rm \frac{x^2-2x-15}{x-3}\div\frac{x-5}{x^2-9}\]Recall your `keep change flip rule` or however you've learned it. \[\large\rm \frac{a}{b}\div\frac{c}{d}\quad=\quad \frac{a}{b}\times\frac{d}{c}\]When we divide by a fraction, we can rewrite it as multiplication of the flip of that second fraction. ~Keep the first fraction the same. ~Change the operation from division to multiplication. ~Flip the second fraction. Do you understand how we can apply this rule to our problem?

jhonyy9 (jhonyy9):

so after used these above wrote by @zepdrix you will can factorizing x^2 -2x -15 = (x-5)(x+3) and x^2 -9=(x-3)(x+3) so after these you will can simplify it easy hope these helped